(v) (7x-2y)/ xy = 5

(8x + 7y)/xy = 15

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Solution: (7x-2y)/ xy = 5 7/y – 2/x = 5…………………………..(i) (8x + 7y)/xy = 15 8/y + 7/x = 15…………………………(ii) Substituting 1/x =m in the given equation we get, – 2m + 7n = 5 => -2 + 7n – 5 = 0 ……..(iii) 7m + 8n = 15 => 7m + 8n – 15 = 0 ……(iv) By cross-multiplication method, we get, m/(-105-(-40)) = n/(-35-30) = 1/(-16-49) m/(-65) = n/(-65) = 1/(-65) m/-65 = 1/-65 m = 1 n/(-65) = 1/(-65) n = 1 m = 1 and n = 1 m = 1/x = 1 n = 1/x = 1 Therefore, x = 1 and y = 1